OpenStudy (anonymous):

Choose the equation of the line passing through the point (-1, -2) and perpendicular to y = 1/4x+ 2.

OpenStudy (anonymous):

y = 4x - 6 y = -4x - 6 y = -4x + 2 y = 4x

OpenStudy (anonymous):

Do you know how to find the slope of a line perpendicular to the another line with a given slope?

OpenStudy (anonymous):

honestly, linear equations are my weakness and everything about them confuses me. I am actually somewhat of a math wiz. I am trying to graduate but I have to retake 9th grade algebra and I dont remember anything about this.

OpenStudy (anonymous):

To find the slope of the line perpendicular to the given line, first find the slope by taking the negative reciprocal of the slope of the given equation. Since the slope of the given equation is the coefficient of x (the number in front of x, 1/4)) what would the negative reciprocal of 1/4 be? hint: flip the fraction and change the sign. What do you get?

OpenStudy (anonymous):

I got it, thanks.

OpenStudy (anonymous):

Did you get -4 for the slope?

OpenStudy (anonymous):

yes I did. is that wrong?

OpenStudy (anonymous):

that's correct - flip 1/4 and change the sign and you end up with a slope of -4!

OpenStudy (anonymous):

y = -4x + 2 ?

OpenStudy (anonymous):

Use the point-slope formula. y-y1=m(x-x1). You know already (because you're so smart) that m (the slope) is -4. Your point is (-1, -2). That means x1=-1 and y1=-2. Plug these in and you get y-(-2)=-4(x-(-1)). Solve that and you have your answer. What do you get. (Hint: -4x +2 may not be right)

OpenStudy (anonymous):

y = -4x - 6 ? i recalculated.

OpenStudy (anonymous):

that's right! Good job

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Good luck with your class - keep asking questions if you need help